OpenStudy (anonymous):

6.In a 45°- 45°- 90° right triangle, the length of the hypotenuse is 15. How long are the legs?

OpenStudy (anonymous):

wait so the other two "legs" are the same right?

OpenStudy (anonymous):

your confusing me

OpenStudy (anonymous):

sure because two angles are the same

OpenStudy (anonymous):

how are you confused?

OpenStudy (anonymous):

I dont understand how to get the numbers or anything, i was out the day my teacher taught this and dont get it at all

OpenStudy (anonymous):



OpenStudy (anonymous):

you have one 90 degree angle, and two 45 degree angles so the legs of the right triangle are equal in length

OpenStudy (anonymous):

so when i divide 90 by sqrt 2 i get 45 sqrt 2. what does that mean?

OpenStudy (anonymous):

i have confused you, forget what i wrote and lets start again from the beginning

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is it ok that both legs of the triangle have the same length?

OpenStudy (anonymous):

because both angles are the same

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then lets call them something. lets call them "a"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now lets find the lenght of the hypotenuse using pythagoras

OpenStudy (anonymous):

we know pythagoras says \(a^2+b^2=c^2\) for a right triangle yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

45^2+45^2=90^2

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

45 45 90 is not the measure of the lengths of the sides, it is the measure of the angles

OpenStudy (anonymous):

pythagoras is about the lengths of the sides, not the message of the angles

OpenStudy (anonymous):

oh...then what do we do for this problem?

OpenStudy (anonymous):

i am getting to that

OpenStudy (anonymous):

s-sorry..

OpenStudy (anonymous):

we start with \[a^2+b^2=c^2\] right? but since the lengths of the legs are both "a" we replace b by a and write \[a^2+a^2=c^2\] ok?

OpenStudy (anonymous):

now we write \(a^2+a^2=2a^2\) so we have \(2a^2=c^2\) and we want to find c

OpenStudy (anonymous):

okay so then it would be 45^2 +45^2=90^2 ? right?

OpenStudy (anonymous):

take the square root to find the length of c. we get \[\sqrt{2a^2}=c\]

OpenStudy (anonymous):

and we know that \[\sqrt{a^2}=a\] so what we really have is \[\sqrt{2}a=c\]

OpenStudy (anonymous):

so 4050= 90^2?

OpenStudy (anonymous):

in other words, in plane english, if i know the lengths of the sides of a 45-45 -90 right triangle, then i can find the hypotenuse by multiplying by \(\sqrt{2}\)

OpenStudy (anonymous):

and if i know the length of the hypotenuse, i can find the length of the sides by dividing by \(\sqrt{2}\)

OpenStudy (anonymous):

im still confused.....this isnt helping me very much

OpenStudy (anonymous):

please don't square the measures of the angles. it is the lengths of the sides ok let me say it as simply as i can your hypotenuse is 15 right? the length of the side is therefore \(\frac{15}{\sqrt{2}}\)

OpenStudy (anonymous):

how do i solve the problem........ i understand that is the length of the side....is that the anwser?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you want a decimal use a calculator and find \(15\div \sqrt{2}\)

OpenStudy (anonymous):

but its asking for how long the legs are... not the sides

OpenStudy (anonymous):

the legs are the sides

OpenStudy (anonymous):

they are synonyms

OpenStudy (anonymous):

so all the legs eqaul 15/ sqrt 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes, both legs have length \(\frac{15}{\sqrt{2}}\)

OpenStudy (anonymous):

if i knew the lenght of the legs was 10, then the length of the hypotenuse would be \(10\sqrt{2}\) and if i know that the length of the hypotenuse is 50 then i know the length of the legs is \(\frac{50}{\sqrt{2}}\)

OpenStudy (anonymous):

in other words, so find the hypotenuse multiply by \(\sqrt{2}\) and do find the legs divide the hypotenuse by \(\sqrt{2}\)

myininaya (myininaya):

 By Mr. Pythagoras and his Pythagorean theorem we have \[a^2+a^2=15^2\] Where a is the length of the legs \[2a^2=15^2\] \[a^2=\frac{15^2}{2}\] \[a=\sqrt{\frac{15^2}{2}}=\frac{\sqrt{15^2}}{\sqrt{2}}=\frac{15}{\sqrt{2}} \] I know this is what sat already said. I just thought maybe I would give a go at explaining it since there still seemed to be confusion.